Sodium metal and chlorine gas can be produced industrially by passing an electric current through molten sodium chloride: 2 NaCl(1)= Na(1)+Cl2(g)
what valume of Cl2 at 293K (20 degrees C) and 1.00 atm can be obtained from 68.0 NaCl using this reaction? (M NaCl= 58.44 g/mol; R= 0.0821 L•atm/[K•mol])
1
Expert's answer
2017-06-21T08:33:06-0400
2 NaCl(1)= 2 Na(1)+Cl2(g)
n (Cl2) = ½ n (NaCl)
n = m/M
n (NaCl) = 68.0/58.44 = 1.2 mol
n (Cl2) = V/22.4
V (Cl2) = n (Cl2) · 22.4 = 1.2 · 22.4 = 26.9 L